# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v1

@@ II-C Signal Recovery Scheme
Due to the DFT computation, our proposed recovery scheme can be classified as a DFT-based recovery technique. Other algorithms under this category are the Fourier-Prony method[7], $B^{2}R^{2}$[8], and LASSO-$B^{2}R^{2}$[9]. The advantage of our proposed recovery scheme is that it does not restrict the signal to decay over time (such as in $B^{2}R^{2}$ and LASSO-$B^{2}R^{2}$) nor does it require the signal to be periodic (such as in Fourier-Prony method). The window function ensures that sharp discontinuities are avoided at the periodic boundaries when the length-$N$ sampled signal is repeated. Consequently, spectral leakage is minimized. The vanishing signal amplitude requirement in $B^{2}R^{2}$ and LASSO-$B^{2}R^{2}$ may require long observation time to estimate the modulo residue samples. The computational complexities of these algorithms grow fast with the observation length. In contrast, the proposed algorithm follows the principle of STFT and the observation length of a segment can be controlled. That is, we can either choose to unfold few overlapping segments with large observation length or many overlapping segments with small observation length. In addition, the sliding DFT approach enables processing of modulo samples in short frames. This is suitable for real-time applications with low latency requirement.
@@ III-A Measuring Recovery Performance
$\mathrm{MSE}=$ $\frac{1}{N}\sum_{n=0}^{N-1}\mathbb{E}\left\{\big{|}\hat{f}[n]-f[n]\big{|}^{2}\right\}$
@@
$=$ $\frac{1}{NI}\sum_{i=1}^{I}\sum_{n=0}^{N-1}\bigg{|}\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]^{2}\right\}_{\mathrm{LPF}}\bigg{|}^{2}$
$+2\mathbb{E}\left\{\epsilon^{(i)}_{\mathrm{LPF}}[n]\right\}\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]^{2}\right\}_{\mathrm{LPF}}$
@@
$=$ $\frac{1}{NI}\sum_{i=1}^{I}\sum_{n=0}^{N-1}\bigg{|}\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]^{2}\right\}_{\mathrm{LPF}}\bigg{|}^{2}$
$+\mathbb{E}\left\{\big{|}\epsilon^{(i)}_{\mathrm{LPF}}[n]\big{|}^{2}\right\}.$ (11)
The second lines comes from partitioning the original signal into $I$ length-$N$ segments. The third line follows from the expression of the unfolded signal in equation (II-C) and from doing some algebraic manipulation. The fourth line comes from the fact that the quantization noise has zero mean. Thus, the second term in the third line can be removed.
@@ III-B Statistical Model for Quantization Noise
Since MSE involves the mean square of the quantization noise, it is crucial to understand the statistical properties of the quantization noise. The statistical nature of the quantization noise comes from the dithered quantization framework established in [32]. This framework is appropriate for the modulo ADC since the folding operation ensures that the signal does not overload the quantizer. The ‘no overloading’ property of the modulo ADC, together with the triangle dither $d[n]$, guarantees that the quantization noise is a white process and its power does not depend on the input, i.e. $\mathbb{E}\left\{|\epsilon[n]|^{2}\big{|}f_{\lambda^{\prime}}[n]\right\}=\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}$ (see [32, Theorem 2]). For triangle dither $d[n]$ described in Section II, the induced quantization noise has power $\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}=\frac{1}{4}\left(\frac{2\lambda}{2^{b}}\right)^{2}=\frac{\lambda^{2}}{2^{2b}}$.
@@ III-D Main Result
One key observation to note in Proposition 1 is that the oversampling factor $\mathrm{OF}$ approaches to unity if $|\mathcal{S}_{i}|$ grows strictly slower with $N$. This is true for finite-energy bandlimited functions due to their time-domain decay property (see [9, 8]). That is, $\exists n_{0}$ such that $f[n]<\lambda^{\prime}$ for all $n>n_{0}$. Consequently, $\exists N_{0}$ such that $|\mathcal{S}_{i}|$ does not change $\forall N>N_{0}$. Thus, the sampling rate can be made closer to the Nyquist rate for finite-energy bandlimited signals at the expense of longer observation window.
One drawback of Proposition 1 is that it depends on the input signal $f(t)$. The following proposition provides a sufficient condition for $\mathrm{OF}$ that depends only on the maximum amplitude $\|f(t)\|_{\infty}$ and not on the whole $f(t)$.
@@
Note that Proposition 2 is a sufficient but not a necessary condition. We remove the dependence on $f(t)$ by bounding $\max_{i}|\mathcal{S}_{i}|$ with [9, Equation 14]. However, this bound can be loose in some cases as demonstrated in [9, Table 1].
@@
Our result points out an error in [10, Theorem 1]. The derivation of [10, Theorem 1] did not take into account the contribution of the $\epsilon^{(i)}[n]$ for $n\in\mathcal{S}_{i}^{c}$, resulting in $b>3$ condition. In our Theorem 1, the second term in equation (16) accounts for the impact of $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ on the recovery performance. Although we are interested in the modulo residue samples located at $\mathcal{S}_{i}$, $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ still contributes to $\hat{z}^{(i)}[n]$ through the computation of $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$. In Section IV, we investigate how the second term of (16) is affected by parameters such as segment length, cardinality of $\mathcal{S}_{i}$, and oversampling factor.
@@
$\mathrm{MSE}=\frac{\|f(t)\|_{\infty}^{2}}{\mathrm{OF}\cdot(2^{b}-2)^{2}\cdot\left(\mathrm{OF}-2\right)^{2}},$ (19)
which exhibits an asymptotic growth rate of $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$. This is similar to the asymptotic growth rate established in [10] which neglected spectral leakage. When $K_{\mathrm{SL}}>0$, the asymptotic growth rate of the MSE becomes $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ only. The $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ growth rate of equation (17) can be maintained if we set $N\geq\mathrm{OF}$. However, using a larger window length increases the computational complexity of the sliding window DFT (see Section III-F).
III-E Comparison with Existing Recovery Methods and Conventional ADCs
We now compare the derived theoretical results with other performance guarantees derived for modulo sampling. One popular recovery procedure is the higher-order differences (HoD) approach developed in [6]. Performance guarantee for this algorithm under a bounded noise setting (e.g., quantization noise) has been analyized in [6, Theorem 6]. More precisely, they showed that for finite-energy bandlimited signals, the noisy unfolded samples can be recovered from noisy modulo samples up to an unknown additive constant, i.e.,
@@ III-E Comparison with Existing Recovery Methods and Conventional ADCs
where $p\in\mathbb{Z}$ is not known, if the sampling rate is at least $2^{\alpha}\pi e\times f_{\mathrm{Nyq}}$. Here, $\alpha\in\mathbb{N}$ is a parameter of the bounded noise and depends on both the maximum amplitude of the noise and the modulo threshold. In contrast, Theorem 1 only requires $\mathrm{OF}\geq 3$ under negligible spectral leakage.
A prediction-based approach [12] showed that $\mathrm{OF}$ can be made arbitrarily close to unity in exchange for a significant increase in prediction filter length. However, it was pointed out in [12, Section III] that their proposed approach “collapses in the presence of quantization noise”. In contrast, our algorithm can work with finite quantization bits. It should also be noted that our proposed algorithm can work at a sampling rate arbitrarily close to the Nyquist rate for finite-energy bandlimited signals and sufficiently large window length $N$ by virtue of Proposition 1.
To demonstrate the advantage of a modulo ADC with 1-bit folding side information over a conventional ADC, we first derive the MSE of a conventional $b$-bit ADC under the non-subtractive dithered quantization framework. For a triangle dither $d[n]\in\left(-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\right)$, the ADC parameter $\lambda$ should be set to $\lambda=\left(1+\frac{1}{2^{b}}\right)\|f(t)\|_{\infty}$. After digital filtering, the quantization noise power (which is also the MSE) becomes
@@
Comparing the MSE guarantee for modulo ADC in Theorem 1 and the derived MSE for a conventional ADC, it can be observed that $\mathrm{MSE}_{\mathrm{mod}}=\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ while $\mathrm{MSE}_{\mathrm{conv}}=\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$ under sufficiently large $b$. The fast decay rate of $\mathrm{MSE}_{\mathrm{mod}}$ with respect to $\mathrm{OF}$ is due to the reduction of the ADC parameter $\lambda$. Smaller ADC range entails smaller quantization bins. This demonstrates the superior performance of modulo ADCs over conventional ADCs in oversampled systems.
@@
III-F Computational Complexity
We now derive the computational complexity of the proposed algorithm in terms of the signal length $N_{0}$, window length $N$, and window roll-off $\alpha$ ($0\leq\alpha\leq 1$). The number of segments depends on the ratio $\frac{N_{0}}{N}$, and the parameter $\alpha$ dictates how large is the overlap between consecutive segments. For large $N_{0}$, the number of segments to be processed by the algorithm is in $\mathcal{O}\left(\frac{N_{0}}{N\left(1-\alpha/2\right)}\right)$. Looking at the complexity of the residual recovery algorithm at each frame, the calculation of $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}$ is the most computationally-expensive operation. More precisely, this calculation is in $\mathcal{O}\left(N^{3}\right)$. Considering all the segments to be processed, the overall computational complexity of the proposed recovery method is in $\mathcal{O}\left(\frac{N_{0}N^{2}}{1-\alpha/2}\right)$.
Based on the derived computational complexity of the proposed algorithm, the speed of the proposed sliding DFT-based recovery scheme can be improved by choosing a small window length. However, a small window length also needs a higher $\alpha$ to compensate for the increased spectral leakage. We also note that the recovery algorithm in [10] is a special case of our proposed recovery algorithm with $N=N_{0}$ and $\alpha=0$, i.e., rectangular window with no overlap. Hence, its complexity is in $\mathcal{O}(N_{0}^{3})$ Since $N$ is typically chosen to be much smaller than $N_{0}$, our proposed sliding DFT-based recovery is $\mathcal{O}\left(\frac{N_{0}^{2}(1-\alpha/2)}{N^{2}}\right)$ faster than the recovery algorithm in [10].

### v2

@@ II-C Signal Recovery Scheme
Due to the DFT computation, our proposed recovery scheme can be classified as a DFT-based recovery technique. Other algorithms under this category are the Fourier-Prony method[10], USLSE[12], $B^{2}R^{2}$[11], and LASSO-$B^{2}R^{2}$[13]. The advantage of our proposed recovery scheme is that it does not restrict the signal to decay over time (such as in $B^{2}R^{2}$ and LASSO-$B^{2}R^{2}$) nor does it require the signal to be periodic (such as in Fourier-Prony method). The window function ensures that sharp discontinuities are avoided at the periodic boundaries when the length-$N$ sampled signal is repeated. Consequently, spectral leakage is minimized. The vanishing signal amplitude requirement in $B^{2}R^{2}$ and LASSO-$B^{2}R^{2}$ may require long observation time to estimate the modulo residue samples. The computational complexities of these algorithms grow fast with the observation length. In contrast, the proposed algorithm follows the principle of STFT and the observation length of a segment can be controlled. That is, we can either choose to unfold few overlapping segments with large observation length or many overlapping segments with small observation length. In addition, the sliding DFT approach enables processing of modulo samples in short frames. This is suitable for real-time applications with low latency requirement.
@@ III-A Measuring Recovery Performance
$\mathrm{MSE}=$ $\frac{1}{N_{0}}\sum_{n=0}^{N_{0}-1}\mathbb{E}\left\{\big{|}\hat{f}[n]-f[n]\big{|}^{2}\right\}$
@@
$=$ $\frac{1}{NI}\sum_{i=1}^{I}\sum_{n=0}^{N-1}\bigg{|}\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]\right\}_{\mathrm{LPF}}\bigg{|}^{2}$
$+2\mathbb{E}\left\{\epsilon^{(i)}_{\mathrm{LPF}}[n]\right\}\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]\right\}_{\mathrm{LPF}}$
@@
$=$ $\frac{1}{NI}\sum_{i=1}^{I}\sum_{n=0}^{N-1}\bigg{|}\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]\right\}_{\mathrm{LPF}}\bigg{|}^{2}$
$\;+\mathbb{E}\left\{\big{|}\epsilon^{(i)}_{\mathrm{LPF}}[n]\big{|}^{2}\right\},$ (11)
where $N_{0}$ is the signal length. The second line comes from partitioning the original signal into $I$ length-$N$ segments. The third line follows from the expression of the unfolded signal in equation (II-C) and from doing some algebraic manipulation. The fourth line comes from the fact that the quantization noise has zero mean. Thus, the second term in the third line can be removed.
@@ III-B Statistical Model for Quantization Noise
Since MSE involves the mean square of the quantization noise, it is crucial to understand the statistical properties of the quantization noise. The statistical nature of the quantization noise comes from the dithered quantization framework established in [49]. This framework is appropriate for the modulo ADC since the folding operation ensures that the signal does not overload the quantizer. The ‘no overloading’ property of the modulo ADC, together with the triangle dither $d[n]$, guarantees that the quantization noise is a white process and its power does not depend on the input, i.e. $\mathbb{E}\left\{|\epsilon[n]|^{2}\big{|}f_{\lambda^{\prime}}[n]\right\}=\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}$ (see [49, Theorem 2]). For triangle dither $d[n]$ described in Section II, the induced quantization noise has power $\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}=\frac{1}{4}\left(\frac{2\lambda}{2^{b}}\right)^{2}=\frac{\lambda^{2}}{2^{2b}}$.
@@ III-D Main Result
One key observation to note in Proposition 1 is that the oversampling factor $\mathrm{OF}$ approaches to unity if $|\mathcal{S}_{i}|$ grows strictly slower with $N$. This is true for finite-energy bandlimited functions due to their time-domain decay property (see [13, 11]). That is, $\exists n_{0}$ such that $f[n]<\lambda^{\prime}$ for all $|n|>n_{0}$. Consequently, $\exists N_{0}$ such that $|\mathcal{S}_{i}|$ does not change $\forall N>N_{0}$. Thus, the sampling rate can be made closer to the Nyquist rate for finite-energy bandlimited signals at the expense of longer observation window. Similar trend is observed in the recovery guarantees established in [15]. That is, for signals that are compact $\lambda$-exceedance with parameter $\rho_{e}$, smaller $\mathrm{OF}$ can be used by increasing the number of observations. This $\lambda$-exceedance property is closely related to the time-domain decay property in [13, 11].
The drawback of Proposition 1 is that it depends on the bandlimited input signal $f(t)$. The following proposition provides a sufficient condition for $\mathrm{OF}$ that depends only on the maximum amplitude $\|f(t)\|_{\infty}$ and not on the whole $f(t)$.
@@
Note that Proposition 2 is a sufficient but not a necessary condition. The expression we used to bound $\max_{i}|\mathcal{S}_{i}|$ from above is identical to the bound of the folding instances derived in [13, Equation 8]. However, their proof assumed that $f(t)$ is a bandlimited signal with negligible truncation error. We show in Appendix B that this bound is also valid for the setup considered in this paper. However, this bound can be loose in some cases as demonstrated in [13, Table 1].
@@
Our result points out an error in [14, Theorem 1]. The derivation of [14, Theorem 1] did not take into account the contribution of the $\epsilon^{(i)}[n]$ for $n\in\mathcal{S}_{i}^{c}$, resulting in $b>3$ condition. In our Theorem 1, the second term in equation (16) accounts for the impact of $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ on the recovery performance. Although we are interested in the modulo residue samples located at $\mathcal{S}_{i}$, $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ still contributes to $\hat{z}^{(i)}[n]$ through the computation of $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$. In Section IV, we investigate how the second term of (16) is affected by parameters such as segment length, cardinality of $\mathcal{S}_{i}$, and oversampling factor.
@@
$\mathrm{MSE}$ $=\frac{\|f(t)\|_{\infty}^{2}}{\mathrm{OF}\cdot(2^{b}-2)^{2}\cdot\left(\mathrm{OF}-2\right)^{2}}$
${\color[rgb]{0,0,0}=\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)}$ (19)
This is similar to the asymptotic growth rate established in [14] which neglected spectral leakage. When $K_{\mathrm{SL}}>0$, the asymptotic growth rate of the MSE becomes $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ only. The $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ growth rate of equation (17) can be maintained if we set $N\geq\mathrm{OF}$. However, using a larger window length increases the computational complexity of the sliding window DFT (see Section III-G).
III-E Comparison with Recovery Guarantees of Existing Methods
We now compare the derived theoretical results with other performance guarantees derived for modulo sampling. One popular recovery procedure is the higher-order differences (HoD) approach developed in [7]. Performance guarantee for this algorithm under a bounded noise setting (e.g., quantization noise) has been analyized in [7, Theorem 6]. More precisely, they showed that for finite-energy bandlimited signals, the noisy unfolded samples can be recovered from noisy modulo samples up to an unknown additive constant, i.e.,
@@ III-E Comparison with Recovery Guarantees of Existing Methods
where $p\in\mathbb{Z}$ is not known, if the sampling rate is at least $2^{\alpha}\pi e\times f_{\mathrm{Nyq}}$. Here, $\alpha\in\mathbb{N}$ is a parameter of the bounded noise and depends on both the maximum amplitude of the noise and the modulo threshold. In contrast, Theorem 1 only requires $\mathrm{OF}>3$ under negligible spectral leakage. We also note that $\mathrm{OF}>3$ closely resembles the OF requirement derived in [50] to uniquely identify periodic bandlimited signals under the modulo-DFT sensing model. However, their model first applies DFT on the sampled signal prior to modulo operation. Moreover, the $\mathrm{OF}>3$ result is reminiscent of the spatial oversampling requirement established for DoA estimation using MIMO arrays with modulo nonlinearities [33].
A prediction-based approach [16] showed that $\mathrm{OF}$ can be made arbitrarily close to unity in exchange for a significant increase in prediction filter length. However, it was pointed out in [16, Section III] that their proposed approach “collapses in the presence of quantization noise”. In contrast, our algorithm can work with finite quantization bits. Moreover, our proposed algorithm can work at a sampling rate arbitrarily close to the Nyquist rate for finite-energy bandlimited signals and sufficiently large window length $N$ by virtue of Proposition 1. We also note that the $\mathrm{OF}$ requirement established in [10] for periodic bandlimited signals is
$\mathrm{OF}\geq\frac{N}{N-2|\mathcal{S}|-2},$
where $|\mathcal{S}|$ is the number of folding instances. Evidently, when $K_{\mathrm{SL}}=0$, the sufficient condition on $\mathrm{OF}$ established in Proposition 1 is better than the $\mathrm{OF}$ requirement in [10]. The advantage of Proposition 1 might be attributed to the availability of the 1-bit folding information in our setup.
Finally, there exists a line of work [28, 29, 31] which uses time-domain thresholding to unfold the output produced by generalized modulo sampling models. Unlike the DFT-based recovery methods, thresholding is done in the time domain so spectral leakage is not an issue. Moreover, thresholding-based techniques have established recovery guarantees [28, 29]. However, the sequential estimation of the folding parameters requires that the folding times are well-separated in the time domain. Such separation cannot be guaranteed under ideal modulo sampling with zero hysterisis as demonstrated in [29, Lemma 1]. Moreover, the MSE bound established in [29, Proposition 1] is inversely proportional to the hysterisis $h$. Under ideal modulo nonlinearities, this $h$ goes to zero and so the MSE under the thresholding-based recovery becomes unbounded.
III-F Comparison with Conventional ADCs
To demonstrate the advantage of a modulo ADC with 1-bit folding side information over a conventional ADC, we first derive the MSE of a conventional $b$-bit ADC under the non-subtractive dithered quantization framework. We compare the modulo ADC and conventional ADC under the same bit budget used for amplitude quantization. For a triangle dither $d[n]\in\left(-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\right)$, the ADC parameter $\lambda$ should be set to $\lambda=\left(1+\frac{1}{2^{b}}\right)\|f(t)\|_{\infty}$. After digital filtering, the quantization noise power (which is also the MSE) becomes
@@ III-F Comparison with Conventional ADCs
Comparing the MSE guarantee for modulo ADC in Theorem 1 and the derived MSE for a conventional ADC, it can be observed that $\mathrm{MSE}_{\mathrm{mod}}=\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ while $\mathrm{MSE}_{\mathrm{conv}}=\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$ under sufficiently large $b$. The fast decay rate of $\mathrm{MSE}_{\mathrm{mod}}$ with respect to $\mathrm{OF}$ is due to the reduction of the ADC parameter $\lambda$. Smaller ADC range entails smaller quantization bins. This demonstrates the superior performance of modulo ADCs over conventional ADCs in oversampled systems. Still, the modulo ADC has a slightly higher hardware complexity due to the folding operation and the 1-bit side information $c[n]$.
III-G Computational Complexity
We now derive the computational complexity of the proposed algorithm in terms of the signal length $N_{0}$, window length $N$, window roll-off $\alpha$ ($0\leq\alpha\leq 1$) and oversamping factor $\mathrm{OF}=\frac{1}{\rho}$. The number of segments depends on the ratio $\frac{N_{0}}{N}$, and the parameter $\alpha$ dictates how large is the overlap between consecutive segments. For large $N_{0}$, the number of segments to be processed by the algorithm is in $\mathcal{O}\left(\frac{N_{0}}{N\left(1-\alpha/2\right)}\right)$. Looking at the complexity of the residual recovery algorithm at each frame, the calculation of $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\in\mathbb{C}^{|\mathcal{S}_{i}|\times K}$ is the most computationally-expensive operation. More precisely, this calculation is in $\mathcal{O}\left(K|\mathcal{S}_{\max}|^{2}\right)$, where $|\mathcal{S}_{\max}|=\max_{i}|\mathcal{S}_{i}|$. Since $K\approx(1-\rho)N$ and $|\mathcal{S}_{\max}|\leq K$ under appropriate choice of $\mathrm{OF}$ and $\lambda^{\prime}$, then the complexity of computing $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}$ is in $\mathcal{O}\left((1-\rho)^{3}N^{3}\right)$. Considering all the segments to be processed, the overall computational complexity of the proposed recovery method is in $\mathcal{O}\left(\frac{N_{0}(1-\rho)^{3}N^{2}}{1-\alpha/2}\right)$. Hence, the computational complexity of our proposed algorithm is linear in the signal length $N_{0}$.
@@ III-G Computational Complexity
Based on the derived computational complexity of the proposed algorithm, the speed of the proposed sliding DFT-based recovery scheme can be improved by choosing a small oversampling factor (high $\rho$) that satisfies the sufficient condition in Theorem 1. However, reducing $\mathrm{OF}$ negatively impacts the MSE as shown in equation (17). The speed of the proposed sliding DFT-based recovery scheme can also be improved by choosing a small window length $N$. However, a small window length also needs a higher $\alpha$ to compensate for the increased spectral leakage. Finally, we note that the recovery algorithm in [14] is a special case of our proposed recovery algorithm with $N=N_{0}$ and $\alpha=0$, i.e., rectangular window with no overlap. Hence, its complexity is in $\mathcal{O}((1-\rho)^{3}N_{0}^{3})$. Since $N$ is typically chosen to be much smaller than $N_{0}$, our proposed sliding DFT-based recovery is $\mathcal{O}\left(\frac{N_{0}^{2}(1-\alpha/2)}{N^{2}}\right)$ faster than the recovery algorithm in [14]. Suppose the signal to be processed has $N_{0}=100,000$ samples. Our proposed algorithm with sliding window length $N=256$ and roll-off parameter $\alpha=0.25$ can unfold the whole signal at roughly $7.48\times 10^{-6}$ times the total time needed by the recovery method in [14] to unfold the whole signal.